OpenStudy (anonymous):

Find the volume of the solid generated by revolving the shaded region about the y-axis. x=2 tan(y/5)

OpenStudy (agent0smith):

Using the washer method, \[V = \pi \int\limits\limits_{a}^{b} \left( 2 \tan \frac{ y }{5 } \right)^2 dy \] You didn't post any limits of integration or a picture, so... I can't help much more than that

OpenStudy (agent0smith):

You can integrate that in a few ways... use the fact that tan^x = 1-sec^x, or use this: http://www.kirupa.com/forum/attachment.php?attachmentid=34138&d=1141841621